# Variation of the centered maximal operator

### v2

@@ 2. Transference
for all $2\leq n\leq N$ and $x_{1}<y_{2}<x_{2}<\dots<x_{N-1}<y_{N}<x_{N}$. Moreover, if $l(I_{1})=A$ and $\alpha_{1}\leq a$, then the first summand in each sum above may be omitted (note that $N\geq 2$ implies $K\geq 2$).
@@
We shall prove the second inequality, the proof of the first one is analogous.
Given $x_{n}$, we take $I_{x_{n}}$ as in the conclusions of Lemma 3.1 with accompanying $1\leq k^{\prime}\leq k^{\prime\prime}\leq K$. Then, for $I_{y_{n}}$ centered at $y_{n}$ and such that $r(I_{y_{n}})=r(I_{x_{n}})$, we have
$Mf(x_{n})-Mf(y_{n})\leq\frac{1}{|I_{x_{n}}|}\int_{I_{x_{n}}}f-\frac{1}{|I_{y_{n}}|}\int_{I_{y_{n}}}f\leq\sum_{k=k^{\prime}}^{k^{\prime\prime}}\frac{\alpha_{k}|I_{k}|}{|I_{x_{n}}|}-\frac{\alpha_{k}|I_{k}|}{|I_{y_{n}}|}.$
Observe that $k^{\prime}=1$ cannot happen if $l(I_{1})=A$ and $\alpha_{1}\leq a$ because then $\alpha_{1}\geq Mf(x_{n})$ implies $Mf(y)\geq Mf(x_{n})$ for all $y\leq x_{n}$; this explains the last sentence of the statement. Next, we notice that the numerical inequality $\frac{1}{\rho_{2}}-\frac{1}{\rho_{2}+\rho_{0}}\leq\frac{1}{\rho_{1}}-\frac{1}{\rho_{1}+\rho_{0}}$ holds for $\rho_{0},\rho_{1},\rho_{2}\in(0,\infty)$ when $\rho_{1}\leq\rho_{2}$. Thus, if either $x_{n}\leq c(I_{k})$ or $c(I_{k})\leq y_{n}$, then
$\frac{\alpha_{k}|I_{k}|}{|I_{x_{n}}|}-\frac{\alpha_{k}|I_{k}|}{|I_{y_{n}}|}\leq\alpha_{k}\Big{|}\frac{|I_{k}|}{|J_{I_{k}}(x_{n})|}-\frac{|I_{k}|}{|J_{I_{k}}(y_{n})|}\Big{|}=\alpha_{k}|T_{I_{k}}(x_{n})-T_{I_{k}}(y_{n})|=\alpha_{k}{\rm Var}_{[y_{n},x_{n}]}(T_{I_{k}}).$
In the remaining case, $y_{n}<c(I_{k})<x_{n}$, we add and subtract $\frac{\alpha_{k}|I_{k}|}{|I^{\prime}|}$ with $I^{\prime}$ such that $c(I^{\prime})=c(I_{k})$ and $r(I^{\prime})=r(I_{y_{n}})=r(I_{x_{n}})$, which gives
$\alpha_{k}\bigg{(}\frac{|I_{k}|}{|I_{x_{n}}|}-\frac{|I_{k}|}{|I^{\prime}|}+\frac{|I_{k}|}{|I^{\prime}|}-\frac{|I_{k}|}{|I_{y_{n}}|}\bigg{)}\leq\alpha_{k}\bigg{(}\Big{|}\frac{|I_{k}|}{|J_{I_{k}}(x_{n})|}-\frac{|I_{k}|}{|J_{I_{k}}(c(I_{k}))|}\Big{|}+\Big{|}\frac{|I_{k}|}{|J_{I_{k}}(c(I_{k}))|}-\frac{|I_{k}|}{|J_{I_{k}}(y_{n})|}\Big{|}\bigg{)}$
and the last quantity does not exceed $\alpha_{k}\big{(}2T_{I_{k}}(c(I_{k}))-T_{I_{k}}(x_{n})-T_{I_{k}}(y_{n})\big{)}=\alpha_{k}{\rm Var}_{[y_{n},x_{n}]}(T_{I_{k}})$. ∎
@@
The same conclusion easily follows when $Mf$ has exactly one substantial local maximum. Assume therefore that $N\geq 2$ and $x_{1}<\dots<x_{N}$ are all substantial local maxima of $Mf$. Since $Mf\geq\frac{a+b}{2}$, $T_{I_{k}}(\pm\infty)=0$, and $Mf$ decreases on $[x_{N},\infty)$, arguing as in the proof of Lemma 3.2 gives
${\rm Var}_{[x_{N},\infty)}(Mf)=Mf(x_{N})-\frac{a+b}{2}\leq\sum_{k=1}^{K}\alpha_{k}{\rm Var}_{[x_{N},\infty)}(T_{I_{k}})-\frac{a+b}{2},$
with the first summand omitted if $l(I_{1})=A$ and $\alpha_{1}\leq a$. If $Mf$ is nondecreasing on $(-\infty,x_{1}]$, then
@@
In this case combining the two estimates above with these from Lemma 3.2 gives (1.2), since
@@
Otherwise, if $x_{1}$ is the first local maximum of $Mf$ but $Mf$ is not nondecreasing on $(-\infty,x_{1}]$, then there is a local minimum $y_{1}\in[A,x_{1})$ such that
@@
The above inequality, which splits into two parts, is a consequence of $(a+b)/2\leq Mf$ (first part), and the argument used in the proof of Lemma 3.2 (second part) with the additional nonnegative term $Mf(y_{1})-(a+b)/2$ added. Again the first summand is not used if $l(I_{1})=A$ and $\alpha_{1}\leq a$.
We now analyze separately three cases depending on the relation between $A$, $l(I_{1})$, $a$, and $\alpha_{1}$. If $l(I_{1})>A$, then we choose the second term from the above minimum and obtain (1.2), since after splitting ${\rm Var}(Mf)$ with respect to $(-\infty,x_{1}]$, $[x_{n},x_{n+1}]$, $n=1,\dots,N-1$, and $[x_{N},\infty)$, we obtain
${\rm Var}(Mf)\leq a-\frac{a+b}{2}+\sum_{k=1}^{K}\alpha_{k}{\rm Var}(T_{I_{k}})-\frac{a+b}{2}=\sum_{k=1}^{K}2\alpha_{k}-b\leq{\rm Var}(f)-a.$
If $l(I_{1})=A$ and $\alpha_{1}\leq a$, then ${\rm Var}(T_{I_{1}})$ is not used. Thus, choosing again the second term, we have
${\rm Var}(Mf)\leq\sum_{k=2}^{K}2\alpha_{k}-b=a-\alpha_{1}+\alpha_{1}+\sum_{k=2}^{K}2\alpha_{k}-b-a\leq{\rm Var}(f)-a.$
Finally, if $l(I_{1})=A$ and $\alpha_{1}>a$, then we choose the first term from the relevant minimum. Now, $x_{1}>r(I_{1})$ because otherwise we would have $Mf(y_{1})>a$. Moreover, $Mf(x_{1})\geq\alpha_{1}$ because $x_{1}$ is the first local maximum, while $Mf(x_{1})=\alpha_{1}$ cannot happen because then $Mf(y)=\alpha_{1}$ for all $y\in[c(I_{1}),x_{1}]$ and again $Mf(y_{1})>a$. Consequently, $l(I_{x_{1}})=l(I_{k^{\prime}})$ for some $k^{\prime}>1$ with $\alpha_{k^{\prime}}>\alpha_{1}>a$, and we have $x_{1}\geq c(I_{k^{\prime}})$. Thus, we obtain
@@
since by $x_{1}\geq c(I_{k^{\prime}})\geq c(I_{1})$ the above sum does not exceed $2(\alpha_{1}+\dots+\alpha_{K})-\alpha_{1}-\alpha_{k^{\prime}}$. ∎
@@ 4. Further comments
First, we note that the constants $1$ and $\frac{1}{2}$ in (1.1) are optimal. Indeed, for $a,b\in[0,\infty)$ with $a\geq b$ and $N\in\mathbb{N}$, take $g_{N}\in{\rm BV}_{a}^{b}(\mathbb{R})$ defined by
$g_{N}\coloneqq a{\mathds{1}_{{(-\infty,-1]}}}+N{\mathds{1}_{{(-1,1)}}}+b{\mathds{1}_{{[1,\infty)}}}.$
Then for $N\geq a$, a simple calculation shows ${\rm Var}(g_{N})=2N-a-b$ and ${\rm Var}(Mg_{N})=2N-a-\frac{a+b}{2}$. Letting $N\to\infty$, we obtain $c(\mathbb{R})\geq 1$ and $C_{a}^{b}(\mathbb{R})\leq\frac{a-b}{2}$.
Next, we show that certain variants of Conjecture $\Delta$ and its discrete counterpart hold for other maximal operators. Set
@@
Since the sampling procedure works in the context of $\overrightarrow{M}$ or $\widetilde{M}$ with only minor adjustments, for all $a,b\in\mathbb{C}$ the associated constants $\overrightarrow{c}(\mathbb{R}),\overrightarrow{C_{a}^{b}}(\mathbb{R}),\widetilde{c}(\mathbb{R}),\widetilde{C_{a}^{b}}(\mathbb{R})$ are not worse than their discrete counterparts. Furthermore, when $\mathbb{K}=\mathbb{Z}$ and $a,b\in[0,\infty)$, it is natural to conjecture that $\overrightarrow{c}(\mathbb{Z})=\widetilde{c}(\mathbb{Z})=1$, $\overrightarrow{C_{a}^{b}}(\mathbb{Z})=\max\{0,b-a\}$, and $\widetilde{C_{a}^{b}}(\mathbb{Z})=|b-a|$. Indeed, a suitable variant of (1.3) follows because for each nonnegative $G\in{\rm BV}_{a}^{b}(\mathbb{Z})$ one has $\overrightarrow{M}G\geq\max\{G,b\}$ and $\widetilde{M}G\geq\max\{G,a,b\}$, while the equality $\overrightarrow{M}G(n)=G(n)$ or $\widetilde{M}G(n)=G(n)$ holds whenever $n$ is a local maximum of $\overrightarrow{M}G$ or $\widetilde{M}G$. Considering $G_{N}\in{\rm BV}_{a}^{b}(\mathbb{Z})$ given by $G_{N}\coloneqq a{\mathds{1}_{{-\mathbb{N}}}}+N{\mathds{1}_{{\{0\}}}}+b{\mathds{1}_{{\mathbb{N}}}}$ and $g_{N}\in{\rm BV}_{a}^{b}(\mathbb{R})$ given by $g_{N}\coloneqq\sum_{n\in\mathbb{Z}}G_{N}(n){\mathds{1}_{{[n-1/2,n+1/2)}}}$, we verify that the postulated constants are optimal.
Corollary 4.1.
@@
In the case of indicator functions our proof of (1.2) and (1.3) is simpler. Indeed, let $f\in{\rm BV}(\mathbb{R})$ be as in Theorem 1.1 with $a,b,\alpha_{k}\in\{0,1\}$. Now, if $x$ is a local maximum of $Mf$, then either $f(x)=0$ or $x$ can be adjusted to $x_{*}=c(I_{k})$ for some $k$. On the other hand, if $y$ is a local minimum, then $f(y)=0$. Thus, we can replace $T_{I_{k}}$ by $M{\mathds{1}_{{I_{k}}}}$, since both functions coincide on $\{c(I_{k})\}\cup(\mathbb{R}\setminus I_{k})$. In particular, for each closed interval $J$ whose endpoints are neighboring extrema $x,y$ we obtain
@@
Finally, notice that the method presented in the proof of Theorem 1.1 has some natural limitations. Indeed, during the proof we always choose $I_{y}$ to be the smallest interval centered at $y$ which contains $I_{x}$. This choice is sometimes very far from optimal. To illustrate this obstacle, let us take $g\in{\rm BV}_{0}^{0}(\mathbb{R})$ given by $g\coloneqq{\mathds{1}_{{[-2,-1]\cup[1,2]}}}+C{\mathds{1}_{{[-2,2]}}}$ with $C=10^{6}$, say. Then $Mg$ has local maxima at $-\frac{3}{2},0,\frac{3}{2}$ with the related values being $C+1,C+\frac{1}{2},C+1$, and local minima at $-\frac{1}{2},\frac{1}{2}$ with the related values being $C+\frac{1}{3},C+\frac{1}{3}$. Thus, ${\rm Var}(Mg)=2C+4-\frac{1}{3}\leq 2C+4={\rm Var}(g)$, as desired. However, for $x=0$ we have $I_{x}=(-2,2)$ so that for $y=\frac{1}{2}$ we should take $I_{y}=(-2,3)$ in Theorem 1.1. Unfortunately, the difference between the corresponding averages is much larger than $Mg(0)-Mg(\frac{1}{2})$, therefore ${\rm Var}(Mg)\leq{\rm Var}(g)$ cannot be verified this way.

### v3

@@ 2. Transference
for all $2\leq n\leq N$ and $x_{1}<y_{2}<x_{2}<\dots<x_{N-1}<y_{N}<x_{N}$. Similarly,
$Mf(x_{N})\leq\sum_{k=1}^{K}\alpha_{k}{\rm Var}_{[x_{N},\infty)}(T_{I_{k}})\quad\text{and}\quad Mf(x_{1})\leq\sum_{k=1}^{K}\alpha_{k}{\rm Var}_{(-\infty,x_{1}]}(T_{I_{k}}).$
If $\alpha_{1}\leq a$, then the first summand in each of the four sums above may be omitted.
@@
We shall estimate $Mf(x_{n})-Mf(y_{n})$ and for $Mf(x_{n-1})-Mf(y_{n})$ the proof is analogous. Given $x_{n}$, we take $I_{x_{n}}$ as in the conclusions of Lemma 3.1 with accompanying $1\leq k^{\prime}\leq k^{\prime\prime}\leq K$. Then, for $I_{y_{n}}$ centered at $y_{n}$ and such that $r(I_{y_{n}})=r(I_{x_{n}})$, we have $I_{x_{n}}\subseteq I_{y_{n}}$ and so
$Mf(x_{n})-Mf(y_{n})\leq\frac{1}{|I_{x_{n}}|}\int_{I_{x_{n}}}f-\frac{1}{|I_{y_{n}}|}\int_{I_{y_{n}}}f\leq\bigg{(}\frac{1}{|I_{x_{n}}|}-\frac{1}{|I_{y_{n}}|}\bigg{)}\int_{I_{x_{n}}}f=\sum_{k=k^{\prime}}^{k^{\prime\prime}}\bigg{(}\frac{\alpha_{k}|I_{k}|}{|I_{x_{n}}|}-\frac{\alpha_{k}|I_{k}|}{|I_{y_{n}}|}\bigg{)}.$
Note that if $\alpha_{1}\leq a$ and $k^{\prime}=1$, then $\alpha_{1}\geq Mf(x_{n})$ by Lemma 3.1, contradicting (3.1). Indeed, we can show $Mf(y)\geq Mf(x_{n})$ for all $y<x_{n}$ by taking $I_{y}$ with $r(I_{y})=r(I_{x_{n}})$. This explains the final sentence of the statement. Next, we notice that the numerical inequality $\frac{1}{\rho_{2}}-\frac{1}{\rho_{2}+\rho_{0}}\leq\frac{1}{\rho_{1}}-\frac{1}{\rho_{1}+\rho_{0}}$ holds for $\rho_{0},\rho_{1},\rho_{2}\in(0,\infty)$ when $\rho_{1}\leq\rho_{2}$. Thus, if either $x_{n}\leq c(I_{k})$ or $c(I_{k})\leq y_{n}$, then
$\frac{\alpha_{k}|I_{k}|}{|I_{x_{n}}|}-\frac{\alpha_{k}|I_{k}|}{|I_{y_{n}}|}\leq\alpha_{k}\bigg{|}\frac{|I_{k}|}{|J_{I_{k}}(x_{n})|}-\frac{|I_{k}|}{|J_{I_{k}}(y_{n})|}\bigg{|}=\alpha_{k}|T_{I_{k}}(x_{n})-T_{I_{k}}(y_{n})|=\alpha_{k}{\rm Var}_{[y_{n},x_{n}]}(T_{I_{k}}).$
In the remaining case, $y_{n}<c(I_{k})<x_{n}$, we add and subtract $\frac{\alpha_{k}|I_{k}|}{|I^{\prime}|}$ with $I^{\prime}$ such that $c(I^{\prime})=c(I_{k})$ and $r(I^{\prime})=r(I_{y_{n}})=r(I_{x_{n}})$. This gives
$\alpha_{k}\bigg{(}\frac{|I_{k}|}{|I_{x_{n}}|}-\frac{|I_{k}|}{|I^{\prime}|}+\frac{|I_{k}|}{|I^{\prime}|}-\frac{|I_{k}|}{|I_{y_{n}}|}\bigg{)}\leq\alpha_{k}\bigg{(}\,\bigg{|}\frac{|I_{k}|}{|J_{I_{k}}(x_{n})|}-\frac{|I_{k}|}{|J_{I_{k}}(c(I_{k}))|}\bigg{|}+\bigg{|}\frac{|I_{k}|}{|J_{I_{k}}(c(I_{k}))|}-\frac{|I_{k}|}{|J_{I_{k}}(y_{n})|}\bigg{|}\,\bigg{)}$
and the last quantity does not exceed $\alpha_{k}\big{(}2T_{I_{k}}(c(I_{k}))-T_{I_{k}}(x_{n})-T_{I_{k}}(y_{n})\big{)}=\alpha_{k}{\rm Var}_{[y_{n},x_{n}]}(T_{I_{k}})$.
Similarly, taking $y_{1}\to-\infty$ and $I_{y_{1}}$ centered at $y$ and such that $r(I_{y})=r(I_{x_{1}})$, we obtain
$Mf(x_{1})=\lim_{y_{1}\to-\infty}\bigg{(}\frac{1}{|I_{x_{1}}|}-\frac{1}{|I_{y_{1}}|}\bigg{)}\int_{I_{x_{1}}}f\leq\lim_{y_{1}\to-\infty}\sum_{k=1}^{K}\alpha_{k}{\rm Var}_{[y_{1},x_{1}]}(T_{I_{k}})=\sum_{k=1}^{K}\alpha_{k}{\rm Var}_{(-\infty,x_{1}]}(T_{I_{k}})$
with the first summand omitted if $\alpha_{1}\leq a$. As before, for $Mf(x_{N})$ the proof is analogous. ∎
@@
Assume therefore that $N\geq 1$ and $x_{1}<\dots<x_{N}$ are all representatives. Since $a\geq b$, we obtain $Mf(x)\geq Mf(\infty)$ for all $x\in\mathbb{R}$ and so $Mf$ decreases on $[x_{N},\infty)$. Therefore, Lemma 3.2 gives
${\rm Var}_{[x_{N},\infty)}(Mf)=Mf(x_{N})-\frac{a+b}{2}\leq\sum_{k=1}^{K}\alpha_{k}{\rm Var}_{[x_{N},\infty)}(T_{I_{k}})-\frac{a+b}{2}$
with the first summand omitted if $\alpha_{1}\leq a$. If $Mf$ is nondecreasing on $(-\infty,x_{1}]$, then similarly
@@
In this case, combining (3.2) with the two estimates above and Lemma 3.2 gives (1.2), since
@@
If $Mf$ is not nondecreasing on $(-\infty,x_{1}]$, then $Mf$ has a local minimum at $y_{1}\in[A,x_{1})$ such that
@@
The above inequality, which splits into two parts, follows from $(a+b)/2\leq Mf$ (first part) and the estimate for $Mf(x_{1})$ from Lemma 3.2 with the additional nonnegative term $22Mf(y_{1})-(a+b)/$ 2 added (second part). Again the first summand is not used if $\alpha_{1}\leq a$. Now, if $\alpha_{1}\leq a$, then choosing the second term from the above minimum and applying (3.2) and Lemma 3.2, we obtain
${\rm Var}(Mf)\leq\sum_{k=2}^{K}2\alpha_{k}-b=(a-\alpha_{1})+\Big{(}\alpha_{1}+\sum_{k=2}^{K-1}2\alpha_{k}+\alpha_{K}\Big{)}+(\alpha_{K}-b)-a\leq{\rm Var}(f)-a$
by the separation condition $\alpha_{k}\alpha_{k+1}=0$. Finally, if $\alpha_{1}>a$, then we choose the first term from the relevant minimum. Now $x_{1}>r(I_{1})$, since otherwise we would have $Mf(y_{1})>a$ by taking $I_{y_{1}}$ with $r(I_{y_{1}})=r(I_{1})$. Also, $Mf(x_{1})<\alpha_{1}$ cannot occur because $Mf(r(I_{1}))\geq\alpha_{1}>Mf(-\infty)$, while $x_{1}$ is the leftmost representative. Similarly, $Mf(x_{1})=\alpha_{1}$ would imply $Mf(y)=\alpha_{1}$ for all $y\in[c(I_{1}),x_{1}]$ due to $x_{1}$ being the leftmost representative, leading to the contradiction $Mf(y_{1})>a$. Thus, $l(I_{x_{1}})=l(I_{k^{\prime}})$ for some $k^{\prime}$ with $\alpha_{k^{\prime}}>\alpha_{1}>a$ and $c(I_{k^{\prime}})\leq x_{1}$ by Lemma 3.1. We then obtain
@@
since by $x_{1}\geq c(I_{k^{\prime}})>c(I_{1})$ the sum above does not exceed $2(\alpha_{1}+\dots+\alpha_{K})-\alpha_{1}-\alpha_{k^{\prime}}$. ∎
@@ 4. Further comments
First, we note that the constants $1$ and $\frac{1}{2}$ in (1.1) are optimal. Indeed, for $a,b\in[0,\infty)$ with $a\geq b$ and $N\in\mathbb{N}$, take $g_{N}\in{\rm BV}_{a}^{b}(\mathbb{R})$ defined by $g_{N}\coloneqq a{\mathds{1}_{{(-\infty,-1)}}}+N{\mathds{1}_{{[-1,1)}}}+b{\mathds{1}_{{[1,\infty)}}}.$ Then for $N\geq a$, a simple calculation shows ${\rm Var}(g_{N})=2N-a-b$ and ${\rm Var}(Mg_{N})=2N-a-\frac{a+b}{2}$. Letting $N\to\infty$, we obtain $C(\mathbb{R})\geq 1$ and $c_{a}^{b}(\mathbb{R})\leq\frac{a-b}{2}$.
Next, we show that certain variants of Conjecture $\Delta$ hold for other maximal operators. Set
@@
Since the sampling procedure works in the context of $\overrightarrow{M}$ or $\widetilde{M}$ with only minor adjustments, for all $a,b\in\mathbb{C}$ the associated constants $\overrightarrow{C}(\mathbb{R}),\overrightarrow{c_{a}^{b}}(\mathbb{R}),\widetilde{C}(\mathbb{R}),\widetilde{c_{a}^{b}}(\mathbb{R})$ are not worse than their discrete counterparts. Furthermore, when $\mathbb{K}=\mathbb{Z}$ and $a,b\in[0,\infty)$, it is natural to conjecture that $\overrightarrow{C}(\mathbb{Z})=\widetilde{C}(\mathbb{Z})=1$, $\overrightarrow{c_{a}^{b}}(\mathbb{Z})=\max\{0,b-a\}$, and $\widetilde{c_{a}^{b}}(\mathbb{Z})=|b-a|$. Indeed, a suitable variant of (1.3) follows because for each nonnegative $G\in{\rm BV}_{a}^{b}(\mathbb{Z})$ one has $\overrightarrow{M}G\geq\max\{G,b\}$ and $\widetilde{M}G\geq\max\{G,a,b\}$, while the equality $\overrightarrow{M}G(n)=G(n)$ or $\widetilde{M}G(n)=G(n)$ holds whenever $\overrightarrow{M}G$ or $\widetilde{M}G$ has a local maximum at $n$. Considering $G_{N}\in{\rm BV}_{a}^{b}(\mathbb{Z})$ given by $G_{N}\coloneqq a{\mathds{1}_{{-\mathbb{N}}}}+N{\mathds{1}_{{\{0\}}}}+b{\mathds{1}_{{\mathbb{N}}}}$ and $g_{N}\in{\rm BV}_{a}^{b}(\mathbb{R})$ given by $g_{N}\coloneqq\sum_{n\in\mathbb{Z}}G_{N}(n){\mathds{1}_{{[n-1/2,n+1/2)}}}$, we verify that the postulated constants are optimal.
Remark 4.1.
@@
In the case of indicator functions our proof of (1.2) and (1.3) is simpler. Indeed, let $f\in{\rm BV}(\mathbb{R})$ be as in Theorem 1.1 with $a,b,\alpha_{k}\in\{0,1\}$. Now, if $Mf$ has a local maximum at $x$, then either $f(x)=0$ or $x$ can be adjusted to $x_{*}=c(I_{k})$ for some $k$. Moreover, if $Mf$ has a local minimum at $y$, then $f(y)=0$. Thus, we can replace $T_{I_{k}}$ by $M{\mathds{1}_{{I_{k}}}}$, since both functions coincide on $\{c(I_{k})\}\cup(\mathbb{R}\setminus I_{k})$. In particular, if $Mf$ is monotone on the closed interval $J$ with endpoints $x,y$ as before, then
@@
Finally, notice that the method presented in the proof of Theorem 1.1 has some natural limitations. Indeed, during the proof we always choose $I_{y}$ to be the smallest interval centered at $y$ which contains $I_{x}$. This choice is sometimes very far from optimal. To illustrate this obstacle, let us take $g\in{\rm BV}_{0}^{0}(\mathbb{R})$ given by $g\coloneqq{\mathds{1}_{{[-2,-1)\cup[1,2)}}}+C{\mathds{1}_{{[-2,2)}}}$ with $C=10^{6}$, say. Then $Mg$ has local maxima at $-\frac{3}{2},0,\frac{3}{2}$ with the related values being $C+1,C+\frac{1}{2},C+1$, and local minima at $-\frac{1}{2},\frac{1}{2}$ with the related values being $C+\frac{1}{3},C+\frac{1}{3}$. Thus, ${\rm Var}(Mg)=2C+4-\frac{1}{3}\leq 2C+4={\rm Var}(g)$, as desired. However, for $x=0$ we have $I_{x}=(-2,2)$ so that for $y=\frac{1}{2}$ we should take $I_{y}=(-2,3)$ in Theorem 1.1. Unfortunately, the difference between the corresponding averages is much larger than $Mg(0)-Mg(\frac{1}{2})$, therefore ${\rm Var}(Mg)\leq{\rm Var}(g)$ cannot be verified this way.